OpenStudy (pottersheep):

The sum of reciprocals of two consecutive integets is 11/56. How would I show this ?? (dont need to solve) thanx

OpenStudy (anonymous):

\[\frac{1}{n}+\frac{1}{n+1}=\frac{11}{56}\]go from there

OpenStudy (pottersheep):

Thank you !

OpenStudy (pottersheep):

are you sure it isnt n + 1/n = 11/56?